# Cancellation properties for exotic $4$-dimensional positive scalar curvature metrics

### v2

@@ 3.3. First part of the proof of Theorem B
Let $M^{d}$ be a closed manifold with a $2$-connected $\theta$-structure $\ell:M\to B$, and let $[f]\in\Gamma^{\theta}(M,\ell)$ be given. Let $N^{n}$ be a further closed manifold. Then the map $(f\times\mathrm{id}_{N})^{*}:\mathcal{R}^{+}(M\times N)\to\mathcal{R}^{*}(M\times N)$ is homotopic to the identity, under each of the following hypotheses.
@@
    $d+n\geq 6$ and the cobordism class of the mapping torus $[T(f)]\in\Omega_{d+1}^{\theta}$ vanishes.
@@
_×_$\_\times\_:\Omega_{d}^{\theta}\times\Omega_{n}^{\vartheta}\to\Omega_{d+n}^{\theta\oplus\theta}$
taking products of representatives. If $N^{n}$ comes equipped with a $\vartheta$-structure, then $[f\times\mathrm{id}_{N}]\in\Gamma(M\times N)$ can be lifted to $\Gamma^{\theta\oplus\vartheta}(M\times N)$, and
$[T(f\times\mathrm{id}_{N})]=[T(f)]\times[N]\in\Omega_{d+n}^{\theta\oplus\vartheta}.$
Since $[T(f)]=0$, we get $[T(f\times\mathrm{id}_{N})]=0$.
@@
[…]
@@
Depending on whether $M$ admits a spin structure or not, let $\theta=\mathrm{Spin}$ or $\theta=\mathrm{SO}$. In both cases, we obtain $\theta$-structures $\ell:M\to B$ which are $2$-connected. Also in both cases, the class of $[f]\in\Gamma(M)$ admits a lift to $\Gamma^{\theta}(M)$ (use Remark 3.3), which we also denote $[f]$. In both cases, we have
$[T(f)]=0\in\Omega_{5}^{\theta};$
@@ 4. The bordism computation
and the elements $g_{+}$, $g_{-}$ are necessarily invertible. Therefore, the functions $\delta_{\pm}:g\mapsto\mathrm{sign}(\det(g_{\pm}))$ are locally constant. We get continuous functions $\delta_{\pm}:\mathrm{Aut}(b)\to\{\pm 1\}=\mathbb{Z}/2$, and by restricting to the maximal compact subgroup which is isomorphic to $O(p)\times O(q)$, one proves that these are group homomorphisms, and that $\delta_{+}(g)\delta_{-}(g)=\det(g)$. It is well-known that
@@
    In §(4.1), we recall the Kervaire semicharacteristic; this will entail a proof of (1.8).
@@
    We conclude the argument in §4.3 by constructing $\varphi$ with the claimed properties. Explicit examples prove that $\varphi$ is surjective, and hence injective by step (2). The claim about $\pi_{1}(\mathrm{MT}\mathrm{Spin}(4))$ will then also follow immediately.
@@ 4.1. The Kervaire semicharacteristic
Let $M$ be a closed manifold of dimension [^4] We are interested in the case $m=1$. $14m+$ 1 and let $\mathbb{F}$ be a field. The Kervaire semicharacteristic $\mathrm{Kerv}_{\mathbb{F}}(M)\in\mathbb{Z}/2$ of $M$ over $\mathbb{F}$ is
$\mathrm{Kerv}_{\mathbb{F}}(X):=\sum_{k=0}^{2m}\dim(H^{k}(M;\mathbb{F}))\in\mathbb{Z}/2.$
@@
$\mathrm{Kerv}_{\mathbb{F}}(T(f))=\dim(\mathrm{Eig}_{1}(H^{2m}(f;\mathbb{F})))\in\mathbb{Z}/2.$
@@
collapses for degree reasons ($H^{p}(S^{1};H^{q}(M;\mathbb{F}))$ is twisted by the action of $\pi_{1}(S^{1})=\mathbb{Z}$ induced by $H^{q}(f;\mathbb{F})$). Therefore
$\kappa(T(f);\mathbb{F})=\dim(E^{0,0}_{2})+\sum_{k=1}^{2m}(\dim(E^{0,k}_{2})+\dim(E^{1,k-1}_{2}))=$ (4.8)
$=\sum_{k=0}^{2m-1}(\dim(E^{0,k}_{2})+\dim(E^{1,k}_{2}))+\dim(E^{0,2m}_{2})\in\mathbb{Z}/2.$
@@
$\kappa(T(f);\mathbb{F})=\dim(E^{0,2m}_{2})\in\mathbb{Z}/2;$
@@
Let $M^{4m}$ be closed an oriented, and let $f\in\mathrm{Diff}^{+}(M)$. Then
@@
This is linear algebra: write $V:=H^{2m}(M;\mathbb{R})$, $b:=I_{M}$ and $g:=H^{2m}(f;\mathbb{R})\in\mathrm{Aut}(V,b)$. Pass to the complexification $V_{\mathbb{C}}$ of $V$ and pick an isomorphism $V_{\mathbb{C}}\cong\mathbb{C}^{n}$ which transforms $b_{\mathbb{C}}$ to the standard bilinear form $(x,y)\mapsto x^{\top}y$; then $g_{\mathbb{C}}$ is represented by a matrix $A\in O(n,\mathbb{C})$. Theorem 9 of [20, §XI.5] describes the possible Jordan normal forms of $A\in O(n,\mathbb{C})$, and from the description given there, the claimed formula follows by an easy case-by-case check. ∎
@@ 4.2. Low–dimensional homotopy groups of Madsen–Tillmann spectra
Since $\mathrm{MT}\theta(d)$ is a Thom spectrum, its homotopy groups can be described in terms of cobordism groups. The following is a straightforward consequence of the classical Pontrjagin–Thom theorem as described in full generality in e.g. [38, Chapter IV §7], [40, §II].
@@
For $k\in\mathbb{Z}$, the group $\pi_{k}(\mathrm{MT}\theta(d))$ is the cobordism group of triples $[M^{d+k},\ell,\varphi]$, where $M^{d+k}$ is a closed manifold, $\ell:M\to B(d)$ is a map and $\varphi:TM\to\ell^{*}V_{d}^{\theta}\oplus\mathbb{R}^{k}$ is a stable bundle isomorphism [^5] More abstractly, $\ell$ and $\varphi$ together are a point in $\mathrm{colim}_{n\to\infty}\mathrm{Bun}(TM\oplus\mathbb{R}^{n},V_{d}^{\theta}\oplus\mathbb{R}^{k+n})$..
@@
of spectra, which is also natural in $\theta$. The connecting map $\mathrm{MT}\theta(n)\to\Sigma\mathrm{MT}\theta(n+1)$ is 00-connected, and the colimit $\mathrm{colim}_{n}\Sigma^{n}\mathrm{MT}\theta(n)$ can be identified with the usual Thom spectrum $M\theta$. It follows that for $k<0$
@@
If $B$ is 00-connected (as it is in the cases $B=B\mathrm{SO}$ and $B=\mathrm{Spin}$ we are interested in), we obtain from (4.11) an exact sequence
[…]
@@
  • The map $\pi_{0}(\mathrm{MT}\theta(n))\to\pi_{-1}(\mathrm{MT}\theta(n+1))=\Omega_{n}^{\theta}$ sends $(M,\ell,\varphi)$ to $(M,\iota_{n}\circ\ell,\mathrm{id}_{\mathbb{R}}\oplus\varphi)$ where $\iota_{n}:B(n)\to B(n+1)$ in the inclusion.
@@
For any field $\mathbb{F}$, the map $[M,f,\varphi]\mapsto\mathrm{Kerv}_{\mathbb{F}}(M)$ gives a well-defined splitting of the maps $\mathbb{Z}/2\to\pi_{0}(\mathrm{MT}\mathrm{Spin}(5)),\,\pi_{0}(\mathrm{MT}\mathrm{SO}(5))$. Hence
$\mathbb{Z}/2\cong\pi_{0}(\mathrm{MT}\mathrm{Spin}(5))\to\mathbb{Z}/2^{2}\cong\pi_{0}(\mathrm{MT}\mathrm{SO}(5)),$
and the map is injective.
@@
of the exact sequence (4.12). The first map takes $[N,\ell,\varphi]$ to
@@
and so maps into $2\mathbb{Z}$. On the other hand $[S^{6},*,\varphi]$, where the stable isomorphism $\varphi:\mathbb{R}^{6}\cong TS^{6}$ comes from the usual $TS^{6}\oplus\mathbb{R}=\mathbb{R}^{7}$ is taken to $2$. This gives the commutative diagram claimed in the lemma.
It is moreover proven in [10, Proposition A.6] that the map $M\mapsto\mathrm{Kerv}_{\mathbb{F}}(M)$ assigning to $M$ its Kervaire semicharacteristic gives a well-defined homomorphism $\pi_{0}(\mathrm{MT}\mathrm{SO}(5))\to\mathbb{Z}/2$, for each field $\mathbb{F}$; the same applies to $\pi_{0}(\mathrm{MT}\mathrm{Spin}(5))$ of course. ∎
@@
[…]
with exact columns. The topmost isomorphisms hold since $B\mathrm{Spin}(5)$ and $B\mathrm{SO}(5)$ are $1$-connected and because $\pi_{1}^{\mathrm{st}}=\mathbb{Z}/2$. The two $\pi_{0}$-groups are finite by Lemma 4.13, hence we may replace the $\mathbb{Z}$’s in the bottom by 00. Using Lemma 4.13 once more, we obtain the commutative diagram
@@ 4.3. Construction of the invariants and their evaluation
The second map in (4.16) is given by the infinite loop space structure. So we should subtract from the map $f:S^{1}\to B\mathrm{Diff}^{\theta}(M)$ which is pointed at $(M,\ell)$ the constant map at this same basepoint which is nothing else that the mapping torus of the identity on $M$; in other words the trivial bundle $M\times S^{1}\to S^{1}$. Therefore, the correct formula for (4.16) is
@@
To define $\varphi_{2}$ and $\varphi_{3}$, we use family index theory. For a smooth bundle $\pi:E\to B$ of oriented $d$-dimensional closed manifolds, we denote by $H^{k}(E\to B;\mathbb{R})\to B$ the real vector bundle whose fibre over $b$ is $H^{k}(M;\mathbb{R})$. This is flat. If $d=4$, the bundle $H^{2}(E\to B;\mathbb{R})$ can be written as a direct sum
@@
so that the (fibrewise) intersection form is positive/negative definite on the summands. This splitting is only natural up to homotopy and the individual summands are not flat. Some linear combinations of these bundles arise as family indices of elliptic operators; for example $\sum_{k=0}^{d}(-1)^{d}H^{k}(E\to B;\mathbb{R})\in KO^{0}(B)$ is the family index of the Euler characteristic operator, and similarly $H^{2,+}(E\to B;\mathbb{C})-H^{2,-}(E\to B;\mathbb{C})\in KU^{0}(B)$ is (when $d=4$) the family index of the signature operator. Note, however, that the signature operator cannot easily be viewed as a real operator, which is why complex $K$-theory appears here.
@@
There are spectrum maps $\chi,\sigma:\mathrm{MT}\mathrm{SO}(4)\to\mathrm{KO}$, such that for each bundle $\pi:E\to B$ of closed oriented $4$-manifolds, the induced maps
$B\stackrel{{\scriptstyle\alpha_{E}}}{{\to}}\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4)\stackrel{{\scriptstyle\Omega^{\infty}\chi,\Omega^{\infty}\sigma}}{{\to}}\Omega^{\infty}\mathrm{KO}\simeq\mathbb{Z}\times BO$
@@
It follows that $\varphi=(\varphi_{1},\chi,\sigma):\pi_{1}(\mathrm{MT}\mathrm{SO}(4))\to\mathbb{Z}/2^{3}$ satisfies (4.3) and that $\mathrm{im}(\pi_{1}(\mathrm{MT}\mathrm{Spin}(4))\to\pi_{1}(\mathrm{MT}\mathrm{SO}(4))$ is contained in $\ker(\varphi_{1})$.
@@
For a closed Riemannian $d$-manifold $M$, let $0\to\mathcal{A}^{0}(M)\stackrel{{\scriptstyle d}}{{\to}}\ldots\stackrel{{\scriptstyle d}}{{\to}}\mathcal{A}^{d}(M)\to 0$ be the de Rham complex (it is crucial for us to use $\mathbb{R}$-valued forms). This is a real elliptic complex and the operator $D=d+d^{*}:\mathcal{A}^{\mathrm{ev}}(M)\to\mathcal{A}^{\mathrm{odd}}$ is elliptic. This construction carries over to the family setting, and (4.19) is the index bundle of $D$. The operator $D$ is “universal” in the sense of [10, §3.2], and in loc.cit., we showed how to reformulate the Atiyah–Singer family index theorem in homotopy–theoretic terms. Namely, there is a universal symbol class $\chi=\sigma_{D}\in KO^{0}(\mathrm{MT}\mathrm{SO}(4))$, so that $\Omega^{\infty}\chi\in KO^{0}(\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4))$ has the property that $(\Omega^{\infty}\chi)\circ\alpha_{E}=\mathrm{ind}(D):B\to\Omega^{\infty}\mathrm{KO}$ (to be precise, in [10] we only covered the case of complex $K$-theory, but [8] also deals with the real case). Writing $\chi$ as a map of spectra yields the claimed property.
For $\sigma$, we use the selfduality complex, a variant of the de Rham complex which exists on oriented $4$-manifolds and plays an important role in gauge theory, see e.g. [39, p. 360 and 446]. For oriented $M^{4}$, the Hodge star $\star:\Lambda^{2}TM\to\Lambda^{2}TM$ is an involution, and we let as usual $\Lambda^{2,\pm}TM$ be its ±$\pm 1$-eigenbundles and let $p_{\pm}:\Lambda^{2}TM\to\Lambda^{2,\pm}TM$ the projection maps. We define $d^{+}:=\sqrt{2}p_{+}d:\mathcal{A}^{1}(M)\to\mathcal{A}^{2,+}(M)$ (the factor of $\sqrt{2}$ is here for cosmetic reasons); the selfduality complex is
$0\to\mathcal{A}^{0}(M)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{A}^{1}(M)\stackrel{{\scriptstyle d^{+}}}{{\to}}\mathcal{A}^{2,+}(M)\to 0$ (4.21)
It is not hard to see that (4.21) is an an elliptic complex, so
@@
The whole construction carries over to families. If $E\to B$ is a bundle of $4$-dimensional oriented manifolds, we get a family of (real) elliptic operators $D^{+}$, and the family index is

### v3

@@ 3.4. First part of the proof of Theorem B
Let $M^{d}$ be a closed manifold with a $2$-connected $\theta$-structure $\ell:M\to B$, and let $f]\in\Gamma^{\theta}(M,\ell)$ be given. Let $N^{n}$ be a further closed manifold. Then the map $(f\times\mathrm{id}_{N})^{*}:\mathcal{R}^{+}(M\times N)\to\mathcal{R}^{*}(M\times N)$ is homotopic to the identity, under each of the following hypotheses.
@@
    $d+n\geq 6$ and the cobordism class of the mapping torus $T(f)]\in\Omega_{d+1}^{\theta}$ vanishes.
@@
$\times\_:\Omega_{d}^{\theta}\times\Omega_{n}^{\vartheta}\to\Omega_{d+n}^{\theta\oplus\vartheta}$
taking products of representatives. If $N^{n}$ comes equipped with a $\vartheta$-structure, then $f\times\mathrm{id}_{N}]\in\Gamma(M\times N)$ can be lifted to $\Gamma^{\theta\oplus\vartheta}(M\times N)$, and
$T(f\times\mathrm{id}_{N})]=[T(f)]\times[N]\in\Omega_{d+n}^{\theta\oplus\vartheta}.$
Since $[T(f)]=0$, we get $T(f\times\mathrm{id}_{N})]=0$.
@@
${S^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f,N)}$${B\mathrm{Diff}^{\theta}(M)\times B\mathrm{Diff}^{\vartheta}(N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{M}\times\alpha_{N}}$${\Omega^{\infty}\mathrm{MT}\theta(d)\times\Omega^{\infty}MT\vartheta(n)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${S^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f\times\mathrm{id}_{N}}$${B\mathrm{Diff}^{\theta\oplus\vartheta}(M\times N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{M\times N}}$${\Omega^{\infty}\mathrm{MT}(\theta\oplus\vartheta)(d+n)}$ (3.20)
@@
Depending on whether $M$ admits a spin structure or not, let $\theta=\mathrm{Spin}$ or $\theta=\mathrm{SO}$. In both cases, we obtain $\theta$-structures $\ell:M\to B$ which are $2$-connected. Also in both cases, the class of $f]\in\Gamma(M)$ admits a lift to $\Gamma^{\theta}(M)$ (use Remark 3.5), which we also denote $[f]$. In both cases, we have
$T(f)]=0\in\Omega_{5}^{\theta};$
@@ 4. The bordism computation
and the elements $_{+}, gg_{-}$ are necessarily invertible. Therefore, the functions $\delta_{\pm}:g\mapsto\mathrm{sign}(\det(g_{\pm}))$ are locally constant. We get continuous functions $\delta_{\pm}:\mathrm{Aut}(b)\to\{\pm 1\}=\mathbb{Z}/2$, and by restricting to the maximal compact subgroup which is isomorphic to $O(p)\times O(q)$, one proves that these are group homomorphisms, and that $\delta_{+}(g)\delta_{-}(g)=\det(g)$. It is well-known that
@@
    In §(4.1), we recall the Kervaire semicharacteristic and prove (1.8).
@@
    We conclude the argument in §4.3 by constructing $\varphi$ with the claimed properties. Explicit examples prove that $\varphi$ is surjective, and hence injective by (2). The claim about $\pi_{1}(\mathrm{MT}\mathrm{Spin}(4))$ will then also follow immediately.
@@ 4.1. The Kervaire semicharacteristic
Let $M$ be a closed manifold of dimension [^5] We are interested in the case $m=1$. $4m+1$ and let $\mathbb{F}$ be a field. The Kervaire semicharacteristic $\mathrm{Kerv}_{\mathbb{F}}(M)\in\mathbb{Z}/2$ of $M$ over $\mathbb{F}$ is
$\mathrm{Kerv}_{\mathbb{F}}(X):=\sum_{k=0}^{2m}\dim(H^{k}(M;\mathbb{F}))\pmod{2}\in\mathbb{Z}/2.$
@@
$\mathrm{Kerv}_{\mathbb{F}}(T(f))=\dim(\mathrm{Eig}_{1}(H^{2m}(f;\mathbb{F})))\pmod{2}\in\mathbb{Z}/2.$
@@
collapses for degree reasons. Here $H^{p}(S^{1};H^{q}(M;\mathbb{F}))$ denotes the cohomology twisted by the action of $\pi_{1}(S^{1})=\mathbb{Z}$ induced by $H^{q}(f;\mathbb{F})$. Therefore
$\mathrm{Kerv}_{\mathbb{F}}(T(f))=\dim(E^{0,0}_{2})+\sum_{k=1}^{2m}(\dim(E^{0,k}_{2})+\dim(E^{1,k-1}_{2}))\pmod{2}$= (4.8)
$\sum_{k=0}^{2m-1}(\dim(E^{0,k}_{2})+\dim(E^{1,k}_{2}))+\dim(E^{0,2m}_{2})\pmod{2}\in\mathbb{Z}/2.$
@@
$\mathrm{Kerv}_{\mathbb{F}}(T(f))=\dim(E^{0,2m}_{2})\pmod{2}\in\mathbb{Z}/2;$
@@
Let $M^{4m}$ be closed and oriented, and let $f\in\mathrm{Diff}^{+}(M)$. Then
@@
This is linear algebra: write $V:=H^{2m}(M;\mathbb{R})$, $b:=I_{M}$ and $g:=H^{2m}(f;\mathbb{R})\in\mathrm{Aut}(V,b)$. Pass to the complexification $V_{\mathbb{C}}$ of $V$ and pick an isomorphism $V_{\mathbb{C}}\cong\mathbb{C}^{n}$ which transforms $b_{\mathbb{C}}$ to the standard bilinear form $(x,y)\mapsto x^{\top}y$; then $g_{\mathbb{C}}$ is represented by a matrix $A\in O(n,\mathbb{C})$. Theorem 9 of [23, §XI.5] describes the possible Jordan normal forms of $A\in O(n,\mathbb{C})$, and from the description given there, the claimed formula follows by an easy case-by-case check. ∎
@@ 4.2. Low–dimensional homotopy groups of Madsen–Tillmann spectra
Since $\mathrm{MT}\theta(d)$ is a Thom spectrum, its homotopy groups can be described in terms of cobordism groups. The following is a straightforward consequence of the classical Pontrjagin–Thom theorem as described in full generality in e.g. [41, Chapter IV §7], [45, §II].
@@
For $k\in\mathbb{Z}$, the group $\pi_{k}(\mathrm{MT}\theta(d))$ is the cobordism group of triples $M^{d+k},\ell,\varphi]$, where $M^{d+k}$ is a closed manifold, $\ell:M\to B(d)$ is a map and $\varphi:TM\to\ell^{*}V_{d}^{\theta}\oplus\mathbb{R}^{k}$ is a stable bundle isomorphism [^6] More abstractly, $\ell$ and $\varphi$ together are a point in $\mathrm{colim}_{n\to\infty}\mathrm{Bun}(TM\oplus\mathbb{R}^{n},V_{d}^{\theta}\oplus\mathbb{R}^{k+n})$..
@@
of spectra, which is natural in $\theta$. The connecting map $\mathrm{MT}\theta(n)\to\Sigma\mathrm{MT}\theta(n+1)$ is 0-connected, and the colimit $\mathrm{colim}_{n}\Sigma^{n}\mathrm{MT}\theta(n)$ can be identified with the usual Thom spectrum $M\theta$. It follows that for $k<0$
@@
If $B$ is 0-connected (as it is in the cases $B=B\mathrm{SO}$ and $B=\mathrm{Spin}$ we are interested in), we obtain from (4.11) an exact sequence
${\pi_{0}(\mathrm{MT}\theta(n+1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{0}(\Sigma^{\infty}B(n+1)_{+})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}\textstyle{\mathbb{Z}}$${\pi_{0}(\mathrm{MT}\theta(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{-1}(\mathrm{MT}\theta(n+1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\Omega_{n}^{\theta}}$${\pi_{-1}(\Sigma^{\infty}B(n+1)_{+})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$ (4.12)
@@
  • The map $\pi_{0}(\mathrm{MT}\theta(n))\to\pi_{-1}(\mathrm{MT}\theta(n+1))=\Omega_{n}^{\theta}$ sends $(M,\ell,\varphi)$ to $(M,\iota_{n}\circ\ell,\mathrm{id}_{\mathbb{R}}\oplus\varphi)$ where $\iota_{n}:B(n)\to B(n+1)$ is the inclusion.
@@
For any field $\mathbb{F}$, the map $M,f,\varphi]\mapsto\mathrm{Kerv}_{\mathbb{F}}(M)$ gives a well-defined splitting of the maps $\mathbb{Z}/2\to\pi_{0}(\mathrm{MT}\mathrm{Spin}(5)),\,\pi_{0}(\mathrm{MT}\mathrm{SO}(5))$. Hence
$\pi_{0}(\mathrm{MT}\mathrm{Spin}(5))\cong\mathbb{Z}/2;\;\pi_{0}(\mathrm{MT}\mathrm{SO}(5))\cong\mathbb{Z}/2^{2},$
and the map $\pi_{0}(\mathrm{MT}\mathrm{Spin}(5))\to\pi_{0}(\mathrm{MT}\mathrm{SO}(5))$ is injective.
@@
of the exact sequence (4.12). The first map takes $N,\ell,\varphi]$ to
@@
and so maps into $2\mathbb{Z}$. On the other hand $S^{6},*,\varphi]$, where the stable isomorphism $\varphi:\mathbb{R}^{6}\cong TS^{6}$ comes from the usual $TS^{6}\oplus\mathbb{R}=\mathbb{R}^{7}$, is taken to $2$. This gives the commutative diagram claimed in the lemma.
It is moreover proven in [13, Proposition A.6] that the map $M\mapsto\mathrm{Kerv}_{\mathbb{F}}(M)$ assigning to $M$ its Kervaire semicharacteristic gives a well-defined homomorphism $\pi_{0}(\mathrm{MT}\mathrm{SO}(5))\to\mathbb{Z}/2$, for each field $\mathbb{F}$; the same applies to $\pi_{0}(\mathrm{MT}\mathrm{Spin}(5))$ of course. ∎
@@
${\pi_{1}(\Sigma^{\infty}B\mathrm{Spin}(5)_{+})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\pi_{1}(\Sigma^{\infty}B\mathrm{SO}(5)_{+})=\mathbb{Z}/2\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathrm{MT}\mathrm{Spin}(4))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{1}(\mathrm{MT}\mathrm{SO}(4))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{0}(\mathrm{MT}\mathrm{Spin}(5))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{0}(\mathrm{MT}\mathrm{SO}(5))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{=}\textstyle{\mathbb{Z}}$
with exact columns. The topmost isomorphisms hold since $B\mathrm{Spin}(5)$ and $B\mathrm{SO}(5)$ are $1$-connected and because $\pi_{1}^{\mathrm{st}}=\mathbb{Z}/2$. The two $\pi_{0}$-groups are finite by Lemma 4.13, hence we may replace the $\mathbb{Z}$’s in the bottom by 0. Using Lemma 4.13 once more, we obtain the commutative diagram
@@ 4.3. Construction of the invariants and their evaluation
The second map in (4.16) is given by the infinite loop space structure. So we must subtract from the map $f:S^{1}\to B\mathrm{Diff}^{\theta}(M)$ which is pointed at $(M,\ell)$ the constant map at this same basepoint which is nothing else that the mapping torus of the identity on $M$; in other words the trivial bundle $M\times S^{1}\to S^{1}$. Therefore, the correct formula for (4.16) is
@@
To define $\varphi_{2}$ and $\varphi_{3}$, we exploit the relation between Madsen–Tillmann spectra and family index theory that we worked out in [13]. For a smooth bundle $\pi:E\to B$ of oriented $d$-dimensional closed manifolds, we denote by $H^{k}(E\to B;\mathbb{R})\to B$ the real vector bundle whose fibre over $b$ is $H^{k}(M;\mathbb{R})$. This is flat. If $d=4$, the bundle $H^{2}(E\to B;\mathbb{R})$ can be written as a direct sum
@@
so that the (fibrewise) intersection form is positive/negative definite on the summands. This splitting is only natural up to homotopy and the individual summands are not flat. Some linear combinations of these bundles arise as family indices of elliptic operators; for example $\sum_{k=0}^{d}(-1)^{d}H^{k}(E\to B;\mathbb{R})\in KO^{0}(B)$ is the family index of the Euler characteristic operator, and similarly $H^{2,+}(E\to B;\mathbb{R})-H^{2,-}(E\to B;\mathbb{R})\in KU^{0}(B)$ is (when $d=4$) the family index of the signature operator.
@@
There are spectrum maps $\chi,\sigma:\mathrm{MT}\mathrm{SO}(4)\to\mathrm{KO}$, such that for each bundle $\pi:E\to B$ of closed oriented $4$-manifolds on finite CW base, the induced maps
$\Omega^{\infty}\chi,\Omega^{\infty}\sigma:B\stackrel{{\scriptstyle\alpha_{E}}}{{\to}}\Omega^{\infty}\mathrm{MT}\mathrm{SO}(4){\to}\Omega^{\infty}\mathrm{KO}\simeq\mathbb{Z}\times BO$
@@
The maps $\chi$ and $\sigma$ induce maps on homotopy groups $\pi_{1}(\mathrm{MT}\mathrm{SO}(4))\to\pi_{1}(\mathrm{KO})=\mathbb{Z}/2$. It follows that $\varphi=(\varphi_{1},\chi_{*},\sigma_{*}):\pi_{1}(\mathrm{MT}\mathrm{SO}(4))\to\mathbb{Z}/2^{3}$ satisfies (4.3) and that $\mathrm{im}(\pi_{1}(\mathrm{MT}\mathrm{Spin}(4))\to\pi_{1}(\mathrm{MT}\mathrm{SO}(4))$ is contained in $\ker(\varphi_{1})$.
@@
For a closed Riemannian $d$-manifold $M$, let $0\to\mathcal{A}^{0}(M)\stackrel{{\scriptstyle d}}{{\to}}\ldots\stackrel{{\scriptstyle d}}{{\to}}\mathcal{A}^{d}(M)\to 0$ be the de Rham complex (we use $\mathbb{R}$-valued forms). This is a real elliptic complex and the operator $D=d+d^{*}:\mathcal{A}^{\mathrm{ev}}(M)\to\mathcal{A}^{\mathrm{odd}}(M)$ is elliptic. This construction carries over to the family setting, and (4.19) is the index bundle of $D$.
To realize (4.20) as the index bundle of a family of real elliptic operators, we use the selfduality complex, a variant of the de Rham complex which exists on oriented $4$-manifolds and plays an important role in gauge theory, see e.g. [42, p. 360 and 446]. For oriented $M^{4}$, the Hodge star $\star:\Lambda^{2}TM\to\Lambda^{2}TM$ is an involution, and we let as usual $\Lambda^{2,\pm}TM$ be its $\pm 1$-eigenbundles and let $p_{\pm}:\Lambda^{2}TM\to\Lambda^{2,\pm}TM$ be the projection maps. We define $d^{+}:=\sqrt{2}p_{+}d:\mathcal{A}^{1}(M)\to\mathcal{A}^{2,+}(M)$ (the factor of $\sqrt{2}$ is here for cosmetic reasons and makes the $D^{+}$ below into a Dirac operator); the selfduality complex is
$0\to\mathcal{A}^{0}(M)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{A}^{1}(M)\stackrel{{\scriptstyle d^{+}}}{{\to}}\mathcal{A}^{2,+}(M)\to 0.$ (4.21)
This is an elliptic complex, so
@@
Again, the construction carries over to families. If $E\to B$ is a bundle of $4$-dimensional oriented manifolds, we get a family of (real) elliptic operators $D^{+}$, and the family index is